Problem 3 After the ice storm, you decide to check on your friend who lives at the top of a hill with an incline of 30 degrees. Due to the ice, your car wont make it up the 105-meter road to her house. With the aid of metal toe cleats, even with ice covering the road, you manage to walk up the hill at 2 When you are 5 meters from the top you lose your toe cleats and start sliding. Assume your mass is 50 kg. (a)[4 pts] What is your acceleration as you slide down the ice-covered road? (b)(4 pts] What is your initial momentum, the instant you slip? (c)[6 pts] How much time does it take you to get back to the bottom of the hill? (d)[5 pts] What is your velocity when you reach the bottom of the hill? (e)(4 pts] What is your momentum the instant you reach the bottom of the hill?
Answer #1

Given,

m = 50 kg ; theta = 30 deg ; D = 105 m ; v = 1/2 m/s ; x = 5 m;

a)Since no friction,

a = g sin(theta)

a = 9.8 sin30 = 4.9 m/s^2

a = 4.9 m/s^2

b)P = m v

P = 0

Hence, P = 0

c)s = ut + 1/2 at^2

s = 105 - 5 = 100

t = sqrt (2 x 100/4.9) = 6.4 s

Hence, t = 6.4 s

d)v = u + at

v = 0 + 4.9 x 6.4 = 31.36 m/s

Hence, v = 31.36 m/s

e)P' = 0